# Graph Signal Processing for Global Stock Market Volatility Forecasting

### v1

@@ 5.3 Experiment Results
v-GSPHAR
vs GSPHAR
Index DM Stats p-Value DM Stats p-Value DM Stats p-Value DM Stats p-Value DM Stats p-Value
FCHI $\mathbf{2.414}$ $\mathbf{0.008}$ $\mathbf{9.383}$ $\mathbf{0}$ $\mathbf{3.332}$ $\mathbf{0}$ $\mathbf{7.360}$ $\mathbf{0}$ $-0.078$ $0.531$
AEX $\mathbf{7.361}$ $\mathbf{0}$ $\mathbf{27.021}$ $\mathbf{0}$ $\mathbf{14.82}$ $\mathbf{0}$ $\mathbf{16.156}$ $\mathbf{0}$ $1.404$ $0.080$
BFX $\mathbf{9.260}$ $\mathbf{0}$ $\mathbf{14.242}$ $\mathbf{0}$ $\mathbf{3.035}$ $\mathbf{0.001}$ $\mathbf{12.484}$ $\mathbf{0}$ $1.153$ $0.124$
STOXX50E $\mathbf{8.186}$ $\mathbf{0}$ $\mathbf{12.208}$ $\mathbf{0}$ $\mathbf{2.777}$ $\mathbf{0.003}$ $\mathbf{20.129}$ $\mathbf{0}$ $\mathbf{4.745}$ $\mathbf{0}$
IBEX $\mathbf{8.216}$ $\mathbf{0}$ $\mathbf{15.957}$ $\mathbf{0}$ $\mathbf{6.455}$ $\mathbf{0}$ $\mathbf{14.270}$ $\mathbf{0}$ $\mathbf{3.547}$ $\mathbf{0}$
GDAXI $\mathbf{10.085}$ $\mathbf{0}$ $\mathbf{37.947}$ $\mathbf{0}$ $\mathbf{6.074}$ $\mathbf{0}$ $\mathbf{22.738}$ $\mathbf{0}$ $\mathbf{7.345}$ $\mathbf{0}$
AORD $\mathbf{4.613}$ $\mathbf{0}$ $\mathbf{3.610}$ $\mathbf{0}$ $0.770$ $0.221$ $\mathbf{8.864}$ $\mathbf{0}$ $\mathbf{2.223}$ $\mathbf{0.013}$
FTSE $\mathbf{11.468}$ $\mathbf{0}$ $\mathbf{2.789}$ $\mathbf{0.003}$ $\mathbf{2.898}$ $\mathbf{0.002}$ $\mathbf{21.631}$ $\mathbf{0}$ $0.481$ $0.315$
MXX $\mathbf{2.034}$ $\mathbf{0.021}$ $\mathbf{6.491}$ $\mathbf{0}$ $\mathbf{2.58}$ $\mathbf{0.005}$ $\mathbf{15.513}$ $\mathbf{0}$ $0.980$ $0.164$
IXIC $\mathbf{3.711}$ $\mathbf{0}$ $\mathbf{6.044}$ $\mathbf{0}$ $\mathbf{6.441}$ $\mathbf{0}$ $\mathbf{16.923}$ $\mathbf{0}$ $\mathbf{6.551}$ $\mathbf{0}$
SSMI $\mathbf{5.491}$ $\mathbf{0}$ $\mathbf{27.432}$ $\mathbf{0}$ $\mathbf{14.34}$ $\mathbf{0}$ $\mathbf{13.679}$ $\mathbf{0}$ $\mathbf{3.621}$ $\mathbf{0}$
SPX $\mathbf{6.003}$ $\mathbf{0}$ $\mathbf{9.312}$ $\mathbf{0}$ $\mathbf{7.732}$ $\mathbf{0}$ $\mathbf{23.259}$ $\mathbf{0}$ $\mathbf{4.375}$ $\mathbf{0}$
RUT $\mathbf{5.559}$ $\mathbf{0}$ $0.779$ $0.218$ $0.083$ $0.467$ $\mathbf{11.084}$ $\mathbf{0}$ $0.979$ $0.164$
DJI $\mathbf{8.924}$ $\mathbf{0}$ $\mathbf{7.925}$ $\mathbf{0}$ $\mathbf{12.742}$ $\mathbf{0}$ $\mathbf{22.485}$ $\mathbf{0}$ $\mathbf{2.825}$ $\mathbf{0.002}$
BSESN $\mathbf{17.734}$ $\mathbf{0}$ $\mathbf{9.389}$ $\mathbf{0}$ $\mathbf{9.085}$ $\mathbf{0}$ $\mathbf{24.455}$ $\mathbf{0}$ $\mathbf{7.845}$ $\mathbf{0}$
NSEI $\mathbf{19.561}$ $\mathbf{0}$ $\mathbf{11.122}$ $\mathbf{0}$ $\mathbf{12.249}$ $\mathbf{0}$ $\mathbf{27.296}$ $\mathbf{0}$ $\mathbf{7.910}$ $\mathbf{0}$
KS11 $\mathbf{7.128}$ $\mathbf{0}$ $\mathbf{12.125}$ $\mathbf{0}$ $\mathbf{7.097}$ $\mathbf{0}$ $\mathbf{19.955}$ $\mathbf{0}$ $\mathbf{10.284}$ $\mathbf{0}$
BVSP $\mathbf{10.374}$ $\mathbf{0}$ $\mathbf{8.513}$ $\mathbf{0}$ $\mathbf{10.490}$ $\mathbf{0}$ $\mathbf{11.808}$ $\mathbf{0}$ $\mathbf{2.104}$ $\mathbf{0.018}$
HSI $\mathbf{5.070}$ $\mathbf{0}$ $\mathbf{13.739}$ $\mathbf{0}$ $\mathbf{2.889}$ $\mathbf{0.002}$ $\mathbf{6.463}$ $\mathbf{0}$ $\mathbf{3.854}$ $\mathbf{0}$
KSE $\mathbf{12.620}$ $\mathbf{0}$ $\mathbf{6.263}$ $\mathbf{0}$ $\mathbf{10.527}$ $\mathbf{0}$ $\mathbf{15.704}$ $\mathbf{0}$ $\mathbf{11.588}$ $\mathbf{0}$
N225 $\mathbf{17.899}$ $\mathbf{0}$ $\mathbf{3.805}$ $\mathbf{0}$ $\mathbf{8.245}$ $\mathbf{0}$ $\mathbf{20.918}$ $\mathbf{0}$ $\mathbf{8.258}$ $\mathbf{0}$
SSEC $\mathbf{14.961}$ $\mathbf{0}$ $\mathbf{6.670}$ $\mathbf{0}$ $\mathbf{14.393}$ $\mathbf{0}$ $\mathbf{34.431}$ $\mathbf{0}$ $\mathbf{14.830}$ $\mathbf{0}$
OSEAX $\mathbf{20.460}$ $\mathbf{0}$ $\mathbf{18.208}$ $\mathbf{0}$ $\mathbf{4.175}$ $\mathbf{0}$ $\mathbf{29.922}$ $\mathbf{0}$ $\mathbf{10.381}$ $\mathbf{0}$
GSPTSE $\mathbf{3.644}$ $\mathbf{0}$ $\mathbf{27.000}$ $\mathbf{0}$ $\mathbf{7.414}$ $\mathbf{0}$ $\mathbf{23.140}$ $\mathbf{0}$ $\mathbf{5.222}$ $\mathbf{0}$
@@
Table 8: The DM test results table of the mid-term forecasting task with bold numbers represents that the GSPHAR is significantly better than the competing model at $5\%$ level.
@@
v-GSPHAR
vs GSPHAR
Index DM Stats p-Value DM Stats p-Value DM Stats p-Value DM Stats p-Value DM Stats p-Value
FCHI $\mathbf{3.776}$ $\mathbf{0}$ $\mathbf{9.002}$ $\mathbf{0}$ $\mathbf{4.033}$ $\mathbf{0}$ $\mathbf{6.758}$ $\mathbf{0}$ $\mathbf{3.836}$ $\mathbf{0}$
AEX $\mathbf{2.858}$ $\mathbf{0.002}$ $\mathbf{7.192}$ $\mathbf{0}$ $\mathbf{4.365}$ $\mathbf{0}$ $\mathbf{6.206}$ $\mathbf{0}$ $\mathbf{3.125}$ $\mathbf{0.001}$
BFX $\mathbf{2.549}$ $\mathbf{0.005}$ $\mathbf{6.445}$ $\mathbf{0}$ $\mathbf{2.885}$ $\mathbf{0.002}$ $\mathbf{4.488}$ $\mathbf{0}$ $\mathbf{2.67}$ $\mathbf{0.004}$
STOXX50E $\mathbf{3.957}$ $\mathbf{0}$ $\mathbf{9.735}$ $\mathbf{0}$ $\mathbf{5.205}$ $\mathbf{0}$ $\mathbf{5.447}$ $\mathbf{0}$ $\mathbf{3.859}$ $\mathbf{0}$
IBEX $\mathbf{4.437}$ $\mathbf{0}$ $\mathbf{11.208}$ $\mathbf{0}$ $\mathbf{7.212}$ $\mathbf{0}$ $\mathbf{6.434}$ $\mathbf{0}$ $\mathbf{4.530}$ $\mathbf{0}$
GDAXI $\mathbf{3.631}$ $\mathbf{0}$ $\mathbf{9.271}$ $\mathbf{0}$ $\mathbf{4.579}$ $\mathbf{0}$ $\mathbf{6.203}$ $\mathbf{0}$ $\mathbf{4.085}$ $\mathbf{0}$
AORD $\mathbf{2.188}$ $\mathbf{0.014}$ $\mathbf{3.778}$ $\mathbf{0}$ $-0.522$ $0.699$ $1.24$ $0.107$ $1.045$ $0.148$
FTSE $\mathbf{3.611}$ $\mathbf{0}$ $\mathbf{7.421}$ $\mathbf{0}$ $\mathbf{3.059}$ $\mathbf{0.001}$ $1.337$ $0.091$ $\mathbf{3.340}$ $\mathbf{0}$
MXX $1.381$ $0.084$ $\mathbf{5.841}$ $\mathbf{0}$ $1.248$ $0.106$ $\mathbf{11.218}$ $\mathbf{0}$ $\mathbf{1.741}$ $\mathbf{0.041}$
IXIC $\mathbf{2.969}$ $\mathbf{0.001}$ $\mathbf{4.505}$ $\mathbf{0}$ $\mathbf{2.844}$ $\mathbf{0.002}$ $1.138$ $0.128$ $\mathbf{2.386}$ $\mathbf{0.009}$
SSMI $\mathbf{2.607}$ $\mathbf{0.005}$ $\mathbf{7.065}$ $\mathbf{0}$ $\mathbf{4.448}$ $\mathbf{0}$ $\mathbf{4.937}$ $\mathbf{0}$ $\mathbf{2.824}$ $\mathbf{0.002}$
SPX $\mathbf{3.897}$ $\mathbf{0}$ $\mathbf{5.530}$ $\mathbf{0}$ $\mathbf{3.805}$ $\mathbf{0}$ $0.269$ $0.394$ $\mathbf{4.414}$ $\mathbf{0}$
RUT $\mathbf{1.900}$ $\mathbf{0.029}$ $\mathbf{3.065}$ $\mathbf{0.001}$ $0.611$ $0.271$ $1.440$ $0.075$ $0.305$ $0.380$
DJI $\mathbf{4.193}$ $\mathbf{0}$ $\mathbf{4.990}$ $\mathbf{0}$ $\mathbf{2.985}$ $\mathbf{0.001}$ $0.665$ $0.253$ $\mathbf{3.736}$ $\mathbf{0}$
BSESN $\mathbf{5.572}$ $\mathbf{0}$ $\mathbf{8.433}$ $\mathbf{0}$ $\mathbf{8.362}$ $\mathbf{0}$ $\mathbf{8.526}$ $\mathbf{0}$ $\mathbf{4.202}$ $\mathbf{0}$
NSEI $\mathbf{6.111}$ $\mathbf{0}$ $\mathbf{9.776}$ $\mathbf{0}$ $\mathbf{9.058}$ $\mathbf{0}$ $\mathbf{9.041}$ $\mathbf{0}$ $\mathbf{4.504}$ $\mathbf{0}$
KS11 $\mathbf{4.403}$ $\mathbf{0}$ $\mathbf{2.834}$ $\mathbf{0.002}$ $\mathbf{4.536}$ $\mathbf{0}$ $\mathbf{6.473}$ $\mathbf{0}$ $\mathbf{3.646}$ $\mathbf{0}$
BVSP $\mathbf{2.941}$ $\mathbf{0.002}$ $\mathbf{5.760}$ $\mathbf{0}$ $\mathbf{2.332}$ $\mathbf{0.010}$ $-0.584$ $0.720$ $\mathbf{2.266}$ $\mathbf{0.012}$
HSI $\mathbf{2.520}$ $\mathbf{0.006}$ $\mathbf{4.747}$ $\mathbf{0}$ $\mathbf{2.139}$ $\mathbf{0.016}$ $\mathbf{4.679}$ $\mathbf{0}$ $\mathbf{1.959}$ $\mathbf{0.025}$
KSE $\mathbf{5.841}$ $\mathbf{0}$ $\mathbf{2.885}$ $\mathbf{0.002}$ $\mathbf{5.342}$ $\mathbf{0}$ $\mathbf{6.234}$ $\mathbf{0}$ $\mathbf{4.063}$ $\mathbf{0}$
N225 $\mathbf{6.927}$ $\mathbf{0}$ $\mathbf{4.755}$ $\mathbf{0}$ $\mathbf{4.412}$ $\mathbf{0}$ $0.736$ $0.231$ $\mathbf{4.988}$ $\mathbf{0}$
SSEC $\mathbf{5.813}$ $\mathbf{0}$ $\mathbf{5.702}$ $\mathbf{0}$ $\mathbf{7.178}$ $\mathbf{0}$ $\mathbf{5.112}$ $\mathbf{0}$ $\mathbf{6.451}$ $\mathbf{0}$
OSEAX $\mathbf{6.456}$ $\mathbf{0}$ $\mathbf{7.916}$ $\mathbf{0}$ $\mathbf{4.659}$ $\mathbf{0}$ $0.728$ $0.233$ $\mathbf{4.569}$ $\mathbf{0}$
GSPTSE $\mathbf{3.541}$ $\mathbf{0}$ $\mathbf{6.612}$ $\mathbf{0}$ $\mathbf{5.534}$ $\mathbf{0}$ $\mathbf{3.195}$ $\mathbf{0.001}$ $\mathbf{3.488}$ $\mathbf{0}$
@@
Table 9: The DM test results table of the long-term forecasting task with bold numbers represents that the GSPHAR is significantly better than the competing model at $5\%$ level.

### v2

@@ 4.3 Experiment Results
Index DM Stats p-Value DM Stats p-Value DM Stats p-Value DM Stats p-Value
FCHI $\mathbf{2.414}$ $\mathbf{0.008}$ $\mathbf{9.383}$ $\mathbf{0}$ $\mathbf{3.332}$ $\mathbf{0}$ $\mathbf{1.808}$ $\mathbf{0.035}$
AEX $\mathbf{7.361}$ $\mathbf{0}$ $\mathbf{27.021}$ $\mathbf{0}$ $\mathbf{14.82}$ $\mathbf{0}$ $\mathbf{5.877}$ $\mathbf{0}$
BFX $\mathbf{9.260}$ $\mathbf{0}$ $\mathbf{14.242}$ $\mathbf{0}$ $\mathbf{3.035}$ $\mathbf{0.001}$ $\mathbf{5.439}$ $\mathbf{0}$
STOXX50E $\mathbf{8.186}$ $\mathbf{0}$ $\mathbf{12.208}$ $\mathbf{0}$ $\mathbf{2.777}$ $\mathbf{0.003}$ $\mathbf{6.839}$ $\mathbf{0}$
IBEX $\mathbf{8.216}$ $\mathbf{0}$ $\mathbf{15.957}$ $\mathbf{0}$ $\mathbf{6.455}$ $\mathbf{0}$ $\mathbf{4.776}$ $\mathbf{0}$
GDAXI $\mathbf{10.085}$ $\mathbf{0}$ $\mathbf{37.947}$ $\mathbf{0}$ $\mathbf{6.074}$ $\mathbf{0}$ $\mathbf{6.184}$ $\mathbf{0}$
AORD $\mathbf{4.613}$ $\mathbf{0}$ $\mathbf{3.610}$ $\mathbf{0}$ $0.770$ $0.221$ $\mathbf{3.804}$ $\mathbf{0}$
FTSE $\mathbf{11.468}$ $\mathbf{0}$ $\mathbf{2.789}$ $\mathbf{0.003}$ $\mathbf{2.898}$ $\mathbf{0.002}$ $\mathbf{11.007}$ $\mathbf{0}$
MXX $\mathbf{2.034}$ $\mathbf{0.021}$ $\mathbf{6.491}$ $\mathbf{0}$ $\mathbf{2.58}$ $\mathbf{0.005}$ $\mathbf{6.161}$ $\mathbf{0}$
IXIC $\mathbf{3.711}$ $\mathbf{0}$ $\mathbf{6.044}$ $\mathbf{0}$ $\mathbf{6.441}$ $\mathbf{0}$ $\mathbf{3.112}$ $\mathbf{0.001}$
SSMI $\mathbf{5.491}$ $\mathbf{0}$ $\mathbf{27.432}$ $\mathbf{0}$ $\mathbf{14.34}$ $\mathbf{0}$ $\mathbf{2.042}$ $\mathbf{0.021}$
SPX $\mathbf{6.003}$ $\mathbf{0}$ $\mathbf{9.312}$ $\mathbf{0}$ $\mathbf{7.732}$ $\mathbf{0}$ $\mathbf{2.264}$ $\mathbf{0.012}$
RUT $\mathbf{5.559}$ $\mathbf{0}$ $0.779$ $0.218$ $0.083$ $0.467$ $\mathbf{2.260}$ $\mathbf{0.012}$
DJI $\mathbf{8.924}$ $\mathbf{0}$ $\mathbf{7.925}$ $\mathbf{0}$ $\mathbf{12.742}$ $\mathbf{0}$ $\mathbf{2.515}$ $\mathbf{0.006}$
BSESN $\mathbf{17.734}$ $\mathbf{0}$ $\mathbf{9.389}$ $\mathbf{0}$ $\mathbf{9.085}$ $\mathbf{0}$ $\mathbf{11.219}$ $\mathbf{0}$
NSEI $\mathbf{19.561}$ $\mathbf{0}$ $\mathbf{11.122}$ $\mathbf{0}$ $\mathbf{12.249}$ $\mathbf{0}$ $\mathbf{13.094}$ $\mathbf{0}$
KS11 $\mathbf{7.128}$ $\mathbf{0}$ $\mathbf{12.125}$ $\mathbf{0}$ $\mathbf{7.097}$ $\mathbf{0}$ $\mathbf{4.611}$ $\mathbf{0}$
BVSP $\mathbf{10.374}$ $\mathbf{0}$ $\mathbf{8.513}$ $\mathbf{0}$ $\mathbf{10.490}$ $\mathbf{0}$ $\mathbf{4.150}$ $\mathbf{0}$
HSI $\mathbf{5.070}$ $\mathbf{0}$ $\mathbf{13.739}$ $\mathbf{0}$ $\mathbf{2.889}$ $\mathbf{0.002}$ $\mathbf{3.393}$ $\mathbf{0}$
KSE $\mathbf{12.620}$ $\mathbf{0}$ $\mathbf{6.263}$ $\mathbf{0}$ $\mathbf{10.527}$ $\mathbf{0}$ $\mathbf{4.666}$ $\mathbf{0}$
N225 $\mathbf{17.899}$ $\mathbf{0}$ $\mathbf{3.805}$ $\mathbf{0}$ $\mathbf{8.245}$ $\mathbf{0}$ $\mathbf{10.018}$ $\mathbf{0}$
SSEC $\mathbf{14.961}$ $\mathbf{0}$ $\mathbf{6.670}$ $\mathbf{0}$ $\mathbf{14.393}$ $\mathbf{0}$ $\mathbf{12.338}$ $\mathbf{0}$
OSEAX $\mathbf{20.460}$ $\mathbf{0}$ $\mathbf{18.208}$ $\mathbf{0}$ $\mathbf{4.175}$ $\mathbf{0}$ $\mathbf{19.028}$ $\mathbf{0}$
GSPTSE $\mathbf{3.644}$ $\mathbf{0}$ $\mathbf{27.000}$ $\mathbf{0}$ $\mathbf{7.414}$ $\mathbf{0}$ $0.506$ $0.306$
@@
Table 8: The DM test results table of the mid-term forecasting task with bold numbers represents that the GSPHAR is significantly better than the competing model at the $5\%$ significance level.
@@
Index DM Stats p-Value DM Stats p-Value DM Stats p-Value DM Stats p-Value
FCHI $\mathbf{3.776}$ $\mathbf{0}$ $\mathbf{9.002}$ $\mathbf{0}$ $\mathbf{4.033}$ $\mathbf{0}$ $\mathbf{4.899}$ $\mathbf{0}$
AEX $\mathbf{2.858}$ $\mathbf{0.002}$ $\mathbf{7.192}$ $\mathbf{0}$ $\mathbf{4.365}$ $\mathbf{0}$ $\mathbf{5.265}$ $\mathbf{0}$
BFX $\mathbf{2.549}$ $\mathbf{0.005}$ $\mathbf{6.445}$ $\mathbf{0}$ $\mathbf{2.885}$ $\mathbf{0.002}$ $\mathbf{4.520}$ $\mathbf{0}$
STOXX50E $\mathbf{3.957}$ $\mathbf{0}$ $\mathbf{9.735}$ $\mathbf{0}$ $\mathbf{5.205}$ $\mathbf{0}$ $\mathbf{6.182}$ $\mathbf{0}$
IBEX $\mathbf{4.437}$ $\mathbf{0}$ $\mathbf{11.208}$ $\mathbf{0}$ $\mathbf{7.212}$ $\mathbf{0}$ $\mathbf{7.853}$ $\mathbf{0}$
GDAXI $\mathbf{3.631}$ $\mathbf{0}$ $\mathbf{9.271}$ $\mathbf{0}$ $\mathbf{4.579}$ $\mathbf{0}$ $\mathbf{5.293}$ $\mathbf{0}$
AORD $\mathbf{2.188}$ $\mathbf{0.014}$ $\mathbf{3.778}$ $\mathbf{0}$ $0.522$ $0.301$ $-1.418$ $0.922$
FTSE $\mathbf{3.611}$ $\mathbf{0}$ $\mathbf{7.421}$ $\mathbf{0}$ $\mathbf{3.059}$ $\mathbf{0.001}$ $\mathbf{2.612}$ $\mathbf{0.005}$
MXX $1.381$ $0.084$ $\mathbf{5.841}$ $\mathbf{0}$ $1.248$ $0.106$ $\mathbf{29.416}$ $\mathbf{0}$
IXIC $\mathbf{2.969}$ $\mathbf{0.001}$ $\mathbf{4.505}$ $\mathbf{0}$ $\mathbf{2.844}$ $\mathbf{0.002}$ $\mathbf{3.356}$ $\mathbf{0}$
SSMI $\mathbf{2.607}$ $\mathbf{0.005}$ $\mathbf{7.065}$ $\mathbf{0}$ $\mathbf{4.448}$ $\mathbf{0}$ $\mathbf{2.171}$ $\mathbf{0.015}$
SPX $\mathbf{3.897}$ $\mathbf{0}$ $\mathbf{5.530}$ $\mathbf{0}$ $\mathbf{3.805}$ $\mathbf{0}$ $-1.358$ $0.913$
RUT $\mathbf{1.900}$ $\mathbf{0.029}$ $\mathbf{3.065}$ $\mathbf{0.001}$ $0.611$ $0.271$ $0.908$ $0.182$
DJI $\mathbf{4.193}$ $\mathbf{0}$ $\mathbf{4.990}$ $\mathbf{0}$ $\mathbf{2.985}$ $\mathbf{0.001}$ $-0.753$ $0.774$
BSESN $\mathbf{5.572}$ $\mathbf{0}$ $\mathbf{8.433}$ $\mathbf{0}$ $\mathbf{8.362}$ $\mathbf{0}$ $\mathbf{6.293}$ $\mathbf{0}$
NSEI $\mathbf{6.111}$ $\mathbf{0}$ $\mathbf{9.776}$ $\mathbf{0}$ $\mathbf{9.058}$ $\mathbf{0}$ $\mathbf{6.872}$ $\mathbf{0}$
KS11 $\mathbf{4.403}$ $\mathbf{0}$ $\mathbf{2.834}$ $\mathbf{0.002}$ $\mathbf{4.536}$ $\mathbf{0}$ $\mathbf{5.460}$ $\mathbf{0}$
BVSP $\mathbf{2.941}$ $\mathbf{0.002}$ $\mathbf{5.760}$ $\mathbf{0}$ $\mathbf{2.332}$ $\mathbf{0.010}$ $-1.801$ $0.964$
HSI $\mathbf{2.520}$ $\mathbf{0.006}$ $\mathbf{4.747}$ $\mathbf{0}$ $\mathbf{2.139}$ $\mathbf{0.016}$ $\mathbf{5.156}$ $\mathbf{0}$
KSE $\mathbf{5.841}$ $\mathbf{0}$ $\mathbf{2.885}$ $\mathbf{0.002}$ $\mathbf{5.342}$ $\mathbf{0}$ $\mathbf{7.625}$ $\mathbf{0}$
N225 $\mathbf{6.927}$ $\mathbf{0}$ $\mathbf{4.755}$ $\mathbf{0}$ $\mathbf{4.412}$ $\mathbf{0}$ $0.300$ $0.382$
SSEC $\mathbf{5.813}$ $\mathbf{0}$ $\mathbf{5.702}$ $\mathbf{0}$ $\mathbf{7.178}$ $\mathbf{0}$ $\mathbf{5.985}$ $\mathbf{0}$
OSEAX $\mathbf{6.456}$ $\mathbf{0}$ $\mathbf{7.916}$ $\mathbf{0}$ $\mathbf{4.659}$ $\mathbf{0}$ $\mathbf{7.765}$ $\mathbf{0}$
GSPTSE $\mathbf{3.541}$ $\mathbf{0}$ $\mathbf{6.612}$ $\mathbf{0}$ $\mathbf{5.534}$ $\mathbf{0}$ $\mathbf{1.704}$ $\mathbf{0.044}$
@@
Table 9: The DM test results table of the long-term forecasting task with bold numbers represents that the GSPHAR is significantly better than the competing model at the $5\%$ significance level.